1) (4p + 1)²= 24
2) (2 -5t)²= 12


Answer by checkley77(12844)

1) (4p + 1)²= 24
4P+1=SQRT24
4P+1=4.899
4P=4.899-1
4P=3.899
P=3.899/4
P=.975 ANSWER.
2) (2 -5t)²= 12
2-5T=SQRT12
2-5T=3.464
-5T=3.464-2
-5T=1.464
T=1.464/-5
T=.2928 ANSWER.